### v1

@@ Abstract
Objective. An Oculomotor Brain-Computer Interface (BCI) records neural activity from regions of the brain involved in planning eye movements and translates this activity into control commands. However, previous successful oculomotor BCI studies primarily used invasive microelectrode implants in non-human primates, whereas non-invasive studies in human subjects relied on electrooculogram (EOG) artefacts in electroencephalogram (EEG) data, rather than saccade planning. This study aimed to demonstrate the feasibility of an oculomotor BCI using a minimally invasive endovascular StentrodeTM device implanted near the supplementary motor area of a patient with Amyotrophic Lateral Sclerosis (ALS).
Approach. Self-paced Visually-Guided and Free-Viewing Saccade Tasks were designed, in which the participant performed saccades in four directions (left, right, up, down) while simultaneously endovascular EEG and eye gaze recordings were collected. Visually guided saccades were cued with visual stimuli, whereas the free-viewing saccades were self-directed without explicit cues.
Main results. While the neural responses of visually-guided saccades overlapped with cue-evoked potentials, free-viewing saccades exhibited saccade-related potentials that began shortly before eye movement, peaked approximately 50 ms after saccade onset, and persisted for around 200 ms. In the frequency domain, these responses appeared as a low-frequency synchronisation below 15 Hz. Classification of ‘fixation vs. saccade’ was robust, achieving a mean area under the receiver operating characteristic curve (AUC) values of 0.88 within sessions and 0.86 between sessions. In contrast, classifying saccade direction proved more challenging, yielding within-session AUC values of 0.67 for four-class decoding and up to 0.75 for the best-performing binary comparisons (left vs. up and left vs. down).
Significance. This proof-of-concept study demonstrates the feasibility of an endovascular oculomotor BCI in an ALS patient, establishing a foundation for future oculomotor BCI studies in human subjects.
@@
Suleman Rasheed1,2,4, James Bennett3,4, Peter E. Yoo3,4, Anthony N Burkitt1,2, David B. Grayden1,2
Keywords: Intravascular, BMI, Saccade, ALS
@@ 1 Introduction
A Brain-Computer Interface (BCI) acquires brain signals, analyses them, and translates them into digital commands that can control an output device, such as a computer cursor or robotic arm. Previous BCI research [1, 2] has shown that it is possible to predict limb movements (hands and feet) by analysing brain signals recorded when users imagine moving their limbs. However, for tasks like computer cursor control, this is an indirect mapping, whereas predicting the direction of intended eye movements and mapping them to control commands may be more natural. BCI systems that rely on decoding eye movements from brain activity are called ‘Oculomotor BCIs’ or ‘Eye-Brain Computer Interfaces’.
@@
An oculomotor BCI differs from conventional motor BCIs in two main ways. First, in terms of task, the user is asked to perform voluntary eye movements instead of limb movements. Second, the corresponding brain activity is captured from the oculomotor system (brain regions responsible for planning and controlling eye movements) instead of the motor cortex. Jia et al. [3] implanted invasive microelectrode arrays in the oculomotor brain regions (specifically, the frontal eye field and the supplementary eye field) of monkeys and successfully predicted the direction of eye movements, achieving 67% accuracy for six-target classification. Tracking eye movements using brain signals offers several advantages over camera-based eye tracking; it does not require users to remain in front of a camera, is visually non-intrusive, avoids issues with occlusion and background illumination, can be integrated with virtual reality headsets [4], and may benefit individuals with eyelid-related conditions.
This study explores the feasibility of predicting eye movements from brain signals in individuals with severe disabilities, such as Amyotrophic Lateral Sclerosis (ALS). Specifically, it analyses brain signals corresponding to rapid eye movements that are called ‘saccades’, which occur when the eyes quickly shift from one fixation point to another. This approach is particularly relevant for cursor control, as even able-bodied individuals often combine eye movements with hand (computer mouse) movements to guide the cursor on a screen. Another novelty of this study is that, while previous research used non-invasive EEG or invasive microelectrode arrays to achieve this goal, this study uses a relatively new brain recording sensor called the StentrodeTM [5], which is a minimally invasive endovascular device developed by Synchron Inc. (New York, USA) that records brain activity from within blood vessels. It offers signal quality comparable to epidural and subdural arrays without requiring open brain surgery, which is typically required for conventional invasive BCI sensors [6].
@@
The experiment protocols of previous oculomotor BCI studies typically implement a visually-guided, a memory-guided or a free-viewing saccade paradigm. In ‘visually guided saccade’ [7, 8], the user maintains fixation at the centre of the screen and makes a saccade when a visual target appears. In a ‘delayed visually-guided saccade’ [9, 10, 11], the user continues fixating on the central dot even after the target appears, and executes the saccade only once the central dot disappears. The advantage of this delay is that it provides separate ‘target’ and ‘go’ cues, allowing for clearer dissociation of neural correlates related to saccade planning and execution. Another variant of this paradigm is the ‘auditory-guided saccade’ [12], in which visual cues are replaced by auditory ones. In the ‘memory-guided saccade paradigm’ [13, 3], the user maintains fixation on a central dot while a target briefly appears and then disappears. The user must remember the target location and execute a saccade towards it only after the central fixation dot disappears. This ‘memory’ period engages attention and working memory, increasing cognitive involvement in saccade planning. Unlike visually-guided and memory-guided saccades, which rely on external cues, the ‘free-viewing saccade’ [14, 15] does not involve such cues. The user freely decides when and in which direction to move their eyes. The advantage of this paradigm is that it eliminates external visual or auditory influences, ensuring that the observed neural correlates are not confounded by cue-related activity.
Existing invasive oculomotor BCI studies to date have been conducted in macaques, with no direct comparable studies in humans. These studies reported classification accuracies of 66.7% [3], 32.2% [13], 90% [9], 80.3% [10], and 43% [11] for saccade direction classification involving 6, 8, 16, 8, and 4 targets, respectively. These studies recorded brain signals using invasive microelectrode arrays that required open brain surgery for implantation. The brain signals were recorded from key oculomotor brain regions, such as the Frontal Eye Field (FEF) [3, 9, 11], Supplementary Eye Field (SEF) [3, 9, 11], Dorsolateral Prefrontal Cortex (DLPFC) [3], Lateral Prefrontal Cortex (LPFC) [10], Lateral Intraparietal Cortex (LIP) [13], Intraparietal Sulcus (IPS) [11] and Superior Parietal Lobule (SPL) [11], all of which are involved in saccade planning. In terms of experimental design, studies using delayed visually-guided saccades [9, 10, 11] incorporated both pre-saccade and post-saccade data for decoding, whereas those employing memory-guided saccades [3, 13] achieved strong performance using only pre-saccade data. The latter was possible due to the additional pre-saccade ‘memory’ period in memory-guided saccades, which is absent in visually-guided saccades.
The key contribution of our study is the investigation of an oculomotor BCI system in a human participant. An endovascular Stentrode device is implanted near the user’s Supplementary Motor Area (SMA), enabling it to record neural activity from adjacent frontal oculomotor regions, including the Frontal Eye Field (FEF) and Supplementary Eye Field (SEF), which have been shown to encode saccade planning information [16, 17, 18]. Furthermore, while other invasive oculomotor BCI studies used intracortical microelectrode arrays, Lee et al. [11] used an epidural ECoG implant in two monkeys, achieving 43% accuracy in classifying saccades in four directions (left, right, up and down), with a contralateral binary classification accuracy of approximately 80%. Given that the Stentrode provides signal quality comparable to that of an ECoG implant [6], this study serves as an interesting benchmark to compare our results.
Several studies also used saccade-related features present in EEG data to predict the directions of saccade in humans. Samadi et al. [19] achieved a mean accuracy of 87% with five subjects in a delayed visually-guided saccade task. Similarly, Kastrati et al. [8] applied deep learning to distinguish horizontal visually-guided saccades, achieving 95% accuracy in a large dataset of 357 subjects. Belkacem et al. [20] investigated the impact of changing the visual angle of the saccade on task performance and reported that larger visual angles resulted in higher classification accuracies. Specifically, for a four-class saccade classification task, visual angles of 5, 10, 20, and 30 yielded accuracies of 66.5%, 73.2%, 81.2%, and 86.8%, respectively. Furthermore, while these studies employed controlled experimental designs, Drouin-Picaro et al. [21] conducted a study where 27 subjects watched videos while their gaze positions were recorded, achieving 73% classification accuracy for saccades in left, right, up, and down directions.
A primary limitation of existing EEG-based oculomotor studies is that they use ocular artefacts (electrooculography, EOG) to infer eye movement information. EOG artefacts [22] are orders of magnitude larger than the underlying neural signals and are most prominent in frontal EEG channels [23]. Although effective in practice, these artefacts do not reflect saccade planning and are therefore not suitable for applications involving patients with eye movement disorders or those in a completely locked-in state [24]. To isolate neural features specifically related to saccade planning, we propose applying EOG artefact removal algorithms [22, 25] before performing classification or reporting performing using only pre-saccade data. However, existing EEG-based oculomotor studies did not report such results. The saccade classification results are reported either on the post-saccade period only or from both pre-saccade and post-sacade intervals. In addition, these studies were conducted with participants without motor impairments, highlighting the need for validation in people with disabilities.
@@
This study aimed to identify the neural correlates of saccadic eye movements and decode their onset and direction using brain signals recorded via an endovascular neural interface device implanted near the supplementary motor area of a patient with ALS. To achieve this, self-paced, Visually-Guided and Free-Viewing Saccade Tasks were designed, and the resulting data were analysed off-line. Neural correlates of saccades in both time and frequency domains were characterised. Classification of ‘saccade’ trials vs. ‘fixation’ trials (saccade onset classification) and prediction of direction of saccadic eye movements (saccade direction classification) were investigated, and the results were compared between the Visually-Guided and Free-Viewing Saccade Tasks.
@@ 2 Methods
2.1 Study Participant and Implant Details
Data were collected from a late-stage ALS patient (male in his 70s) with complete paralysis of the extremities. The participant was implanted with an endovascular stent-electrode array (StentrodeTM, Synchron, NY, USA) as part of Synchron’s SWITCH clinical trial (NCT03834857), and was referred to as ‘P4’ in an earlier study [26]. The Stentrode consisted of sixteen platinum electrodes, each with a diameter of 500 µm and spaced, on average, 3 mm apart, mounted on a self-expanding nitinol scaffold measuring 8 $\times$ 40 mm [27, 26, 28]. The device was deployed through the right jugular vein and placed in the superior sagittal sinus, immediately adjacent to the precentral gyrus. The Stentrode was connected via a flexible transvascular lead to a subcutaneous Implantable Receiver-Transmitter Unit (IRTU) located on the right side of the participant’s chest. For additional details on the implantation procedure, see [26, 27].
Data for the Visually-Guided Saccade Task were collected in one session, and the Free-Viewing Saccade Task were recorded in two subsequent sessions. The participant was unable to speak, but retained oculomotor control and was able to use an eye tracker for computer access. Data were collected through a study approved by St. Vincent’s Hospital, Melbourne (SAGE ID: 2022 / PID04863-2020 / SVH03352).
@@ 2.2 Experiment Protocol
Two saccade experiments were designed to study Visually-Guided and Free-Viewing saccades. Both tasks were self-paced, and the user was instructed to wait briefly before performing each saccade to extend the planning interval. The user was asked to make saccades towards four different targets on the screen while data from both the eye tracker and the Stentrode were recorded simultaneously. Eye gaze information was used to label saccade onsets. Figure 1 shows the Graphical User Interface (GUI) for both Visually-Guided and Free-Viewing Tasks.
@@ 2.3 Data Collection
Data recordings were made on a 15” Surface Book 3 with a display resolution of 3240 x 2160 pixels. Neural activity was captured by the Stentrode, while eye movements were recorded using a Tobii PCEye eye tracker. The eye tracker was attached to the bottom of the Surface Book screen, with the user positioned at arm’s length (approximately 50 cm) from the screen. The Tobii Gaze Point API was used to control mouse movements via the eye tracker. The Stentrode data were recorded at 2000 Hz, while gaze positions were recorded at approximately 30 Hz for the Visually-Guided Task and 60 Hz for the Free-Viewing Task. The experiment was programmed in PsychoPy [29] and the Lab Streaming Layer (LSL) [30] was used to synchronise the data across these three different modalities.
@@ 2.4.1 Labelling the Saccade Onsets
The first step in the data pre-processing pipeline was to label the onset and offset of saccadic eye movements in each trial. Initially, a heuristic-based approach was used that labelled the saccade onsets based on the cursor’s distance from the initial point A for each saccade between points A and B. This was followed by a visual inspection of each trial to further correct the timings of saccade onset. Trials with incomplete saccades that failed to complete a bidirectional saccade task within the maximum allowed time (10 s for the Visually-Guided and 6.5 s for the Free-Viewing Task) or those with abnormal cursor trajectories, primarily influenced by the presence of blinks, were excluded from subsequent analyses.
@@ 2.4.2 Extracting Fixation and Saccade Trials
As neural activities across different channels of the Stentrode recordings were highly correlated, Independent Component Analysis (ICA) [32] was applied to separate activity arising from distinct neural sources. Before applying ICA, the data were visually inspected, and two noisy channels were removed from the 15 active channels. The remaining 13 channels were then used to estimate the Independent Components (ICs). A 1 Hz high-pass zero-phase FIR filter was applied to preprocess the data before computing the ICA unmixing matrix.
For the Free-Viewing Task, two data sessions were recorded, and the ICs calculated for each of these sessions may not maintain a one-to-one correspondence. To match and align the corresponding ICs, a group ICA approach was used. Specifically, data from both sessions were combined to compute a single ICA unmixing matrix, which was then applied separately to each session. Although alternative approaches, such as visualising power spectra or topographical maps of individual ICs or applying clustering techniques [33], could have been used, we restricted our analysis to group ICA. This decision was made because, unlike conventional EEG topographical maps, the spatial maps derived from the Stentrode are not yet well characterised in the literature, and such an investigation was beyond the scope of the current work.

### v3

@@ Abstract
Objective. An Oculomotor Brain-Computer Interface (BCI) records neural activity from brain regions involved in planning eye movements and translates this activity into control commands. While previous successful studies have relied on invasive implants in non-human primates or electrooculography (EOG) artefacts in human electroencephalogram (EEG) data, this study aimed to demonstrate the feasibility of an oculomotor BCI using a minimally invasive endovascular StentrodeTM device implanted near the supplementary motor area of a patient with Amyotrophic Lateral Sclerosis (ALS).
Approach. One participant performed self-paced visually-guided and free-viewing saccade tasks in four directions (left, right, up, down) while endovascular EEG and eye gaze recordings were collected. Visually-guided saccades were cued with visual stimuli, whereas free-viewing saccades were self-directed without explicit cues. Brain signals were pre-processed to remove cardiac artefacts, downsampled, and classified using a Random Forest algorithm. For saccade onset classification (fixation vs. saccade), features in time and frequency domains were extracted after xDAWN denoising, while for saccade direction classification, the downsampled time series were classified directly without explicit feature extraction.
Main results. The neural responses of visually-guided saccades overlapped with cue-evoked potentials, while free-viewing saccades exhibited saccade-related potentials that began shortly before eye movement, peaked approximately 50 ms after saccade onset, and persisted for around 200 ms. In the frequency domain, these responses appeared as a low-frequency synchronisation below 15 Hz. Saccade onset classification was robust, achieving mean area under the receiver operating characteristic curve (AUC) scores of 0.88 within sessions and 0.86 across sessions. Saccade direction decoding yielded within-session AUC scores of 0.67 for four-class decoding and up to 0.75 for the best performing binary comparisons (left vs. up and left vs. down).
Significance. This proof-of-concept study demonstrates the feasibility of an endovascular oculomotor BCI in a patient with ALS, establishing a foundation for future oculomotor BCI studies in human subjects.
: J. Neural Eng.
@@
Suleman Rasheed1,2,3, James Bennett3,4, Peter E. Yoo3,4, Anthony N. Burkitt1,2, David B. Grayden1,2
Keywords: Oculomotor BCI, BMI, Intravascular, Endovascular, Stentrode, ALS, Saccade
@@ 1 Introduction
A Brain-Computer Interface (BCI) acquires brain signals, analyses them, and translates them into commands that can control an output device, such as a computer cursor or robotic arm. Previous BCI research [1, 2] has shown that it is possible to predict limb movements (hands and feet) by analysing brain signals recorded when users imagine moving their limbs. However, for tasks like computer cursor control, this is an indirect mapping, whereas predicting the direction of intended eye movements and mapping them to control commands may be more natural. BCI systems that rely on decoding eye movements from brain activity are called ‘Oculomotor BCIs’ or ‘Eye-Brain Computer Interfaces’.
@@
In an oculomotor BCI system, the user is asked to perform voluntary eye movements while brain signals are captured over oculomotor brain regions that contribute to planning and execution of eye movements. For example, Jia et al. [3] implanted invasive microelectrode arrays in the frontal eye field and supplementary eye field of macaques, and achieved 67% accuracy in predicting the direction of eye movements in a six-target classification task. Tracking eye movements using brain signals also offers several advantages over camera-based eye tracking: it does not require users to remain in front of a camera, it is visually non-intrusive, it avoids issues with occlusion and background illumination, it can be integrated with virtual reality headsets [4], and it may benefit individuals with eyelid-related conditions.
Existing oculomotor studies focused on decoding ‘saccades’, which are rapid eye movements that occur when the eyes quickly shift from one fixation point to another. Both invasive and non-invasive studies have demonstrated the capability of predicting saccade direction from brain signals. The invasive oculomotor studies implanted microelectrode arrays in macaques and showed that the direction of saccades can be predicted with high accuracies of 66.7% (for 6 targets) [3], 32.2% (8) [5], 90% (16) [6], 80.3% (8) [7], and 43% (4) [8]. The brain signals were recorded from key oculomotor brain regions, such as the Frontal Eye Field (FEF) [3, 6, 8], Supplementary Eye Field (SEF) [3, 6, 8], Dorsolateral Prefrontal Cortex (DLPFC) [3], Lateral Prefrontal Cortex (LPFC) [7], Lateral Intraparietal Cortex (LIP) [5], Intraparietal Sulcus (IPS) [8], and Superior Parietal Lobule (SPL) [8], In addition to intracortical microelectrodes, Lee et al. [8] implanted an epidural electrocorticography (ECoG) array in macaques and reported classification accuracies of 43% (for 4 targets) and 80% (2) for saccade direction classification, respectively. Similarly, non-invasive studies collected EEG data from able-bodied human subjects and predicted saccade directions with accuracies of 87% (for 8 targets) [9], 95% (2)  [10], and 73% (4) [11].
@@
The experiment protocols of these studies typically implemented visually-guided, memory-guided or free-viewing saccade paradigms. In visually-guided saccades [12, 10], the user maintains fixation at the centre of the screen and makes a saccade when a visual target appears. In delayed visually-guided saccades [6, 7, 8], the user continues fixating on the central dot even after the target appears, and executes the saccade only once the central dot disappears. The advantage of this delay is that it provides separate ‘target’ and ‘go’ cues, allowing for clearer dissociation of neural correlates related to saccade planning and execution. Another variant of this paradigm is auditory-guided saccades [13], in which visual cues are replaced by auditory ones. In the memory-guided saccade paradigm [5, 3], the user maintains fixation on a central dot while a target briefly appears and then disappears. The user must remember the target location and execute a saccade towards it only after the central fixation dot disappears. This ‘memory’ period engages attention and working memory, increasing cognitive involvement in saccade planning. Unlike visually-guided and memory-guided saccades, which rely on external cues, free-viewing saccades [14, 15] do not involve such cues; the user freely decides when and in which direction to move their eyes. The advantage of this paradigm is that it eliminates external visual or auditory influences, ensuring that the observed neural correlates are not confounded by cue-related activity.
Despite the high decoding performance, existing oculomotor studies have key limitations. Invasive studies were conducted exclusively in non-human primates, as they required electrode implantation via craniotomy, which poses clinical safety concerns for use in humans. Consequently, there are no equivalent invasive oculomotor BCI studies in humans. In contrast, non-invasive human studies have typically relied on EOG artefacts present in EEG recordings to infer eye movement information. These artefacts originate from the corneo-retinal dipole, formed by the positively charged cornea and negatively charged retina. Eye movements cause changes in this potential, producing signals that are captured by frontal EEG electrodes. These EOG artefacts confound the underlying neural activity related to saccade planning. Existing EEG oculomotor studies [16, 9, 11, 12, 17, 10] did not apply EOG correction algorithms and included data from post-saccade intervals for decoding, making it difficult to determine whether classification performance reflected true neural signals or EOG artefacts. These limitations highlight the need to validate the neural correlates of saccades in human subjects, without EOG artefacts, and explore their potential for oculomotor BCI applications.
@@
The key contribution of our study is the investigation of an oculomotor BCI system in a human participant with ALS, in whom an endovascular neural interface was implanted without the need for a craniotomy. The Stentrode TM [18] is an endovascular neural interface developed by Synchron Inc. (New York, USA) that records brain activity from within blood vessels. It offers signal quality comparable to epidural and subdural arrays without requiring open brain surgery [19]. The study aimed to identify the neural correlates of saccadic eye movements and decode their onset and direction using brain signals recorded via an endovascular Stentrode implanted near the supplementary motor area. To achieve this, self-paced, Visually-Guided and Free-Viewing Saccade Tasks were designed, and the resulting data were analysed offline. Neural correlates of saccades were characterised in both the time and frequency domains. Classification of ‘saccade’ versus ‘fixation’ trials (onset classification) and prediction of saccade direction (direction classification) were investigated, and the results were compared between the Visually-Guided and Free-Viewing Tasks.
@@ 2 Methods
2.1 Participant and Implant Details
Data were collected from a late-stage ALS patient (male in his 70s) with complete paralysis of the extremities. The participant was unable to speak, but retained oculomotor control and was able to use an eye tracker for computer access. The participant was implanted with an endovascular stent-electrode array (StentrodeTM, Synchron, NY, USA) as part of Synchron’s SWITCH clinical trial (NCT03834857), and was referred to as ‘P4’ in an earlier study [20]. The Stentrode was implanted in the superior sagittal sinus, adjacent to the precentral gyrus [20], near the Supplementary Motor Area (SMA) of the subject. As shown in Figure 1, the electrodes are anatomically proximal to frontal oculomotor regions, particularly the Supplementary Eye Field (SEF) and the Frontal Eye Field (FEF). As these regions are involved in the planning and execution of saccades [21, 22, 23], the recorded signals may capture saccade-related neural features.
Figure 1: Anatomical placement of the Stentrode implant relative to major cortical landmarks in a T1-weighted MRI reconstruction of the participant’s brain. The electrodes are positioned near frontal oculomotor regions, including the Supplementary Eye Field (SEF) and the Frontal Eye Field (FEF), both of which are involved in the planning and execution of saccadic eye movements. The brain anatomy and electrode positions are accurately mapped from the participant’s imaging data, whereas the SEF and FEF regions are approximate and shown for illustrative purposes based on typical anatomical landmarks.
The Stentrode consisted of 16 platinum electrodes, each with a diameter of 500 µm and spaced, on average, 3 mm apart, mounted on a self-expanding nitinol scaffold measuring 8 $\times$ 40 mm [24, 20, 25]. The device was deployed through the right jugular vein and placed in the superior sagittal sinus, immediately adjacent to the precentral gyrus. The Stentrode was connected via a flexible transvascular lead to a subcutaneous Implantable Receiver-Transmitter Unit (IRTU) located on the right side of the participant’s chest. For additional details on the implantation procedure, see [20, 24].
@@ 2.2 Experiment Protocol
Two saccade experiments were designed to study Visually-Guided and Free-Viewing saccades. This comparison was intended to reveal differences between externally cued and self-paced voluntary saccades, respectively. Both tasks were self-paced, and the user was instructed to wait briefly before performing each saccade to extend the planning interval. The user was asked to make saccades towards four different targets on the screen while data from both the eye tracker and the Stentrode were recorded simultaneously. Eye gaze information was used to label saccade onsets. Figure 2 shows the Graphical User Interface (GUI) for both Visually-Guided and Free-Viewing Tasks.
@@ 2.3 Data Collection
Data were collected through a study approved by St. Vincent’s Hospital, Melbourne (SAGE ID: 2022 / PID04863-2020 / SVH03352). Data for the Visually-Guided Task were collected in one session and for the Free-Viewing Task in two subsequent sessions. Data recordings were made on a 15” Surface Book 3 with a display resolution of 3240 x 2160 pixels. Neural activity was captured by the Stentrode, while eye movements were recorded using a Tobii PCEye eye tracker. The eye tracker was attached to the bottom of the Surface Book screen, with the user positioned at arm’s length (approximately 50 cm) from the screen. The Tobii Gaze Point API was used to control mouse movements via the eye tracker. The Stentrode data were recorded at 2000 Hz, while gaze positions were recorded at approximately 30 Hz for the Visually-Guided Task and 60 Hz for the Free-Viewing Task. The experiment was programmed in PsychoPy [26] and the Lab Streaming Layer (LSL) [27] was used to synchronise the data across these three different modalities.
@@ 2.4.1 Labelling the Saccade Onsets
The gaze position data recorded by the eye tracker was used to label the onset and offset of saccadic eye movements in each trial. Initially, a heuristic-based approach was used that labelled the saccade onsets based on the cursor’s distance from the initial point A for each saccade between points A and B. This was followed by a visual inspection of each trial to further correct the timings of saccade onset. Trials with incomplete saccades that failed to complete a bidirectional saccade task within the maximum allowed time (10 s for the Visually-Guided and 6.5 s for the Free-Viewing Task) or those with abnormal cursor trajectories, primarily influenced by the presence of blinks, were excluded from subsequent analyses.
@@ 2.4.2 Extracting Fixation and Saccade Trials
As neural activity across different channels of the original Stentrode recordings was highly correlated, Independent Component Analysis (ICA) [29] was applied to separate neural signals arising from distinct neural sources. This also helped to effectively separate cardiac artefacts from the neural signals. ICA is a blind source separation method that decomposes sensor-level data into statistically independent components, which we refer to as the source space. The original electrode recordings are referred to as the sensor space. ICA assumes the observed brain signals $\boldsymbol{X}(t)\in\mathbb{R}^{n\times T}$, where $n$ is the number of channels and $T$ is the number of time points, are linear mixtures of statistically independent sources $\boldsymbol{S}(t)$, combined through an unknown mixing matrix $\boldsymbol{A}$, such that $\boldsymbol{X}(t)=\boldsymbol{A}\cdot\boldsymbol{S}(t)$. An unmixing matrix $\boldsymbol{W}\approx\boldsymbol{A}^{-1}$ is estimated and applied to the sensor space data to obtain the independent components in source space: $\hat{\boldsymbol{S}}(t)=\boldsymbol{W}\cdot\boldsymbol{X}(t)$. In this study, ICA was implemented using the MNE-Python library (version 1.2.1) [30], employing the FastICA algorithm [31], which estimates $\boldsymbol{W}$ by maximising non-gaussianity of the recovered components under unit-variance constraints. Before applying ICA, the data were visually inspected, and two noisy channels were removed from the 15 active channels. The remaining 13 channels were then used to estimate the Independent Components (ICs). A 1 Hz high-pass zero-phase FIR filter was applied to preprocess the data before computing the ICA unmixing matrix.
For the Free-Viewing Task, two data sessions were recorded, and the ICs calculated for each of these sessions may not maintain a one-to-one correspondence. To align the corresponding ICs across sessions, a group-level ICA approach [32, 33] was used, in which data from both sessions were concatenated to compute a single ICA unmixing matrix. This matrix was then applied separately to each session. Alternative approaches, such as visualising power spectra or topographical maps of individual ICs or applying clustering techniques [34], have been used in literature to establish IC correspondence in EEG data. However, because the topographical maps derived from the Stentrode are not yet well characterised in the literature, such an investigation was beyond the scope of the present study.